40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 7 08:00:11 2016, Steve, Update, SUS, SRM -PRM sat. amp swap SRM.jpgSRM-UR_OK.png
    Reply  Mon Nov 7 16:05:55 2016, gautam, Update, SUS, PRM Sat. Box. Debugging D961289-B2.pdfPRMSatBoxtest.png
       Reply  Tue Nov 8 11:54:38 2016, gautam, Update, SUS, PRM Sat. Box. looks to be fixed 
Message ID: 12606     Entry time: Tue Nov 8 11:54:38 2016     In reply to: 12602
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM Sat. Box. looks to be fixed 

Looks like the PRM Sat. Box is now okay, no evidence of the kind of glitchy behaviour we are used to seeing in any of the 5 channels.

Quote:
 
GV Edit 2230 hrs 7Nov2016: The signs from the last 6 hours has been good - see the attached minute trend plot. Usually, the glitches tend to show up in this sort of time frame. I am not quite ready to call the problem solved just yet, but I have restored the connections to the SRM suspension (the PRM and SRM Sat. Boxes are still switched). I've also briefly checked the SRM alignment, and am able to lock the DRMI, but the lock doesn't hold for more than a few seconds. I am leaving further investigations for tomorrow, let's see how the Sat. Box does overnight.

 

ELOG V3.1.3-